OpenStudy (anonymous):

determine the constant that should be added to the binomial so that it becomes a perfect sqr trinomial. Then write the factorization X^2-x

OpenStudy (anonymous):

make the constant \[\left( \frac{1}{2} \right)^2=\frac{1}{4}\]to get the perfect square trinomial expression\[x^2-x+\frac{1}{4}\]

OpenStudy (anonymous):

so it looks as if the constant can really be anything